# Towards Interpretable and Efficient Attention: Compressing All by Contracting a Few

### v2

@@ 3.2 Contract-and-Broadcast Self-Attention
3.3 Variants of CBSA
In this subsection, we demonstrate that specifying different sets of representatives or their extraction strategies derives variants of CBSA that can be categorized as existing attention mechanisms. In these variants, the initialization, extraction, contraction, and broadcast steps of CBSA may not be explicitly observed, as some steps are simplified or fused together due to their specific number and structure of the representatives.
Agent Attention [30] as CBSA variant. As illustrated in Fig. 3, if we remove the contraction step of CBSA, it derives the Agent Attention, which was proposed to be a generalized form of linear attention. [^9] Since query, key, and value coincide in CBSA (8), the broadcast matrix in Agent Attention can be obtained directly from the extraction (i.e., cross-attention) step. Actually, we argue that, in this case, the contraction step is not removed but fused into the extraction step, since most of the extraction’s purpose already achieved by pooling the input tokens.
We previously extract representatives satisfying the relaxed constraints in (5) via MSCA with an initial guess, yielding an approximate numerical solution to the constraints of the original problem (3). Beyond this, closed-form optimal solutions  [^10] That is, the representatives that not only have the same coding rate as the input tokens but also possess an explicit coefficient matrix allowing them to be linearly expressed by the input tokens. also exist, which can either improve interpretability and efficiency or enhance the expressive capacity of the resulting mechanism.
Softmax attention as CBSA variant. Apparently, the input tokens themselves satisfy the constraints and thus can be directly used as the representatives, i.e., $\boldsymbol{Q}=\boldsymbol{Z},m=N$. In this case, we say the input tokens are self-expressed [33], since data samples are linearly represented over a dictionary composed of themselves. Then we can take a trivial solution for the regularization term where all coefficient matrices reduces to identity matrices. Substituting them into (8) yields the following operator, known as MSSA in white-box transformers [16]:
@@ 3.3 Variants of CBSA
Linear attention as CBSA variant. To analyze the case of orthogonal representatives, we start with the canonical choice: the principal directions of input tokens, along which projections yield the principal components. Specifically, We perform singular value decomposition (SVD) on the token projection within each subspace:
@@
where $\boldsymbol{L}_{k}\in\mathcal{O}(p)$, $\boldsymbol{R}_{k}\in\mathcal{O}(N,p)$, and $\boldsymbol{\Sigma}_{k}$ is a $p\times p$ diagonal matrix of singular values. The columns of $\boldsymbol{L}_{k}$ are known as the principal directions. Then, by right multiplying both sides by $\boldsymbol{R}_{k}$, we have:
$\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}\boldsymbol{R}_{k}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\boldsymbol{R}_{k}^{\top}\boldsymbol{R}_{k}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k},\ \ \forall k\in[K],$ (11)
By comparing (10) with the constraints in (5), we let $\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k},\boldsymbol{A}_{k}=\boldsymbol{R}_{k}$. Substituting them into (7) leads to the following operation:
@@
where $F$ is a function defined on the spectrum of a positive semi-definite matrix which applies $f(\lambda_{i})=\nicefrac{{\epsilon^{2}}}{{(\epsilon^{2}+\lambda_{i})}}$ to the eigenvalues $\{\lambda_{i}\}_{i=1}^{p}$ of the covariance matrix of the token projections. [^11] As $f$ is monotonically decreasing, (12) preserves the principle directions (i.e., representatives) with large variance while suppressing those with vanishing variance [27]. In Appendix B, we prove that this result holds for any set of orthogonal representatives. We find that (12) conceptually resembles linear attention, as it also factorizes the $N\times N$ attention matrix and multiplies the key and value first to linearize the computational complexity. [^12] In fact, linear attention is more expressive than (12), since $(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})^{\top}$ is symmetric, whereas $(\boldsymbol{W}^{v}\boldsymbol{Z})(\boldsymbol{W}^{k}\boldsymbol{Z})^{\top}$ is generally not. How to interpret the transition from such symmetric attention to their asymmetric counterparts remains largely unexplored. For now, one can view it as overparameterization.
@@
Channel attention as CBSA variant. If we assume that the subspace bases $\boldsymbol{U}_{k}$ always correspond to the principal directions of any input token set, which is practically impossible but simplifies the operations, the representatives can be fixed as $\boldsymbol{U}_{k}$ (i.e., $\boldsymbol{Q}=[\boldsymbol{U}_{1}^{\top}\ldots\boldsymbol{U}_{K}^{\top}]^{\top}$), and are thus orthogonal and input-agnostic. Then (12) is reduced to:
@@
$\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{D}_{k}\boldsymbol{U}_{k}^{\top}\boldsymbol{Z},\ \boldsymbol{D}_{k}\doteq\operatorname{Diag}\left([f\big((\boldsymbol{u}_{ki}^{\top}\boldsymbol{Z})(\boldsymbol{u}_{ki}^{\top}\boldsymbol{Z})^{\top}\big)]_{i=1}^{p}\right),$ (13)
@@
which corresponds to TSSA [21] without exploiting token-to-subspace membership. In (13), the feature channels are adaptively scaled based on the second moment of the token projections, whereas typical channel attention mechanisms employ an MLP to predict the scaling factors.
@@
Besides a graph perspective, the above findings can also be interpreted from a dictionary learning perspective, by treating the representatives as atoms of a dictionary. When the representatives are orthogonal and fixed, they form a complete dictionary; when they are orthogonal but input-dependent, they resemble a submatrix of an overcomplete dictionary in compressed sensing [34]. When the input tokens themselves serve as representatives, they constitute a self-expressive dictionary [33].
Figure 3: Overview of different attention mechanisms. Zoom in for better view.
@@ 4 Experiments
In this section, we evaluate the interpretability and efficiency of the proposed inherently interpretable attention, CBSA, and the CBT models built upon it. Since natural images often lie on low-dimensional subspaces [35, 36], we focus on classical visual tasks such as image classification and semantic segmentation, where higher resolutions generally lead to better accuracy [37, 38]. Experimental results copied from their original work will be in grey color, otherwise the results will be reproduced by us in varied settings for fair comparisons.
Baseline and training configuration. We primarily compare CBSA with previous interpretable attention mechanisms grounded in MCR2, along with pioneering and classical Transformer-based models. Table 2 summarizes these baselines with brief descriptions. Results directly copied from the original papers are shown in grey, while the others are reproduced by us under varied settings for fair comparisons. Unless otherwise specified, the training configuration follows that of the baselines, with detailed information provided in Appendix C.
Implementation detail. We over-parameterize the projection step that maps tokens back to the ambient space, where the projection matrix should theoretically be $\boldsymbol{U}_{[K]}$, by a set of independently learned parameter matrices. This trick is also adopted by MSSA [16] and TSSA [21], and its impact has been investigated in [39]. In short, although this relaxation compromises the theoretical rigor, it is crucial for achieving better accuracy. In addition, the positive scalar $\kappa$ in (6) is implemented as a learnable parameter without constraining its sign. This allows the model to flexibly choose between compression and de-compression. We provide the PyTorch implementation in Appendix D.
Table 2: Summary of baselines. Note that these methods are not necessarily limited to the task listed behind, our description here only reflects their usages in this paper.
Method Attention mechanism Complexity Interpretable Task
@@
DEPICT [31] MSSA / MSCA (cross-attention) quadratic / linear semantic segmentation
@@ 4.1 Advantages enabled by interpretability
In this subsection, we show some superior advantages of CBSA over non-interpretable attention mechanisms. The most essential one is its interpretability, which further induces other desirable properties such as robustness and emergent segmentation.
@@
Figure 4: Compact and structured representation. There are 10 classes in total and each class has 200 samples.
Learning compact and structured representation. Similar to MCR2, our optimization objective (3) also aims to learn compact and structured representations, i.e., compressing input tokens towards low-dimensional structures such as a mixture of subspaces. To verify that CBSA can achieve this goal, we evaluate the linear-attention-like variant of CBSA (12) on synthetic data. As illustrated in Fig. 4, we emulate image tokens as points in a 3-dimensional space, where ground-truth clustering memberships are indicated by color, and apply (12) iteratively. Since the operation is directly performed in the ambient space and no parameterized subspace is required, this procedure can be executed in a purely forward manner under supervision.
@@
Figure 5: We measure the compression term of (2) across layers in different models.
@@
Compressing all by contracting a few. To verify our interpretation that CBSA compresses all tokens through the contraction of a few representatives, we conduct two-fold experiments showing that the input tokens are indeed compressed and that this compression is driven by the contraction of the representatives. In Fig. 5, we measure the compression term of (2) using the coding rate (1) and its normalized variant  [^13] All input tokens are first normalized to unit vectors before calculating the normalized coding rate.. The normalized coding rate is invariant to in-place scaling and therefore captures only the angular compression. We observe two pieces of evidence supporting the compression interpretation: 1) representation compactness in the last layer, measured by the coding rate in Fig. 5, positively correlates with classification accuracy in Table 3; 2) in nearly all models, the latter half of the layers exhibit consecutive compression. We also identify a de-compression phenomenon in the earlier layers, which we regard as a “known unknown” requiring further investigation. In Fig. 6, we measure the “compressed volume” of input tokens and their representatives separately after being processed by CBSA (without projecting back to the ambient space, so the results are reported per subspace). We observe that the compressed volumes of the two exhibit highly similar trends across most subspaces.
@@
Figure 6: We measure the compressed volume of input tokens and their representatives using the reduced coding rate across all heads of a model. Results from CBT-T/S are presented here, those of CBT-B/L can be found in Fig. 10. Note that we set $\kappa=1$ to exclude the de-compression cases observed in Fig. 5.
@@
Figure 7: [CLS] attention map. We use $\boldsymbol{A}_{k}\boldsymbol{A}_{k}^{\top}\in\mathbb{R}^{N\times N}$ to empirically approximate the full attention matrix. Based on this, we extract the attention map of the [CLS] token from the early, middle, and late layers of CBT, CRATE, and their hybrid model.
@@
Emergent segmentation properties. Owing to MSSA, segmentation properties emerge in CRATE with merely standard supervised classification training [41]. In Fig. 7, we visualize the attention map of the [CLS] token across different models. We observe that CBT attends to more semantically meaningful regions in the early layers compared to CRATE, but its segmentation properties do not persist in subsequent layers. To address this, we construct a hybrid model, termed CRATE+CBT, where the first half of the attention layers employ MSSA and the latter half employ CBSA. In this hybrid model, segmentation properties not only emerge in the very first layer, but are also progressively enhanced in the following layers, rather than fading as in CBT. Qualitative results supporting this conclusion can be found in Appendix C.
@@
Figure 8: Semantic segmentation on ADE20K. $\dagger$ indicates that CBSA implemented rigorously, i.e., without the overparameterization trick. All results above are evaluated on the ADE20K validation set. Leftmost panel: The dotted horizontal line marks the best result. Right two panels: Random Gaussian noise with zero mean and standard deviation $\sigma$ is independently added to each parameter of the attention layers in the decoder.
@@
Robustness under parameter perturbation. As CBSA models low-dimensional subspaces via their attention heads, with parameter matrices corresponding to subspace bases, perturbing these basis vectors with relatively small noise does not significantly alter the subspaces they span [31]. Consequently, CBSA exhibits extreme robustness to parameter perturbations, whereas the black-box method Segmenter [40] collapses under the same perturbations, as illustrated in the right two panels of Fig. 8.
@@
Table 3: Image classification on ImageNet-1K. The models are trained on images of resolution $224\times 224$ with a patch size of $16$. Note that the CBT models and ViT employ convolutional embedding layers, whereas the CRATE models use linear embeddings.
@@
Table 4: Fair comparisons on ImageNet-1K. All models employ convolutional embedding layers and ISTA feedforward blocks. Their only difference lies in the attention mechanism. Specially, the Agent models do not follow the original implementation in [30]; instead, they are implemented as CBSA without the contraction step, as derived in Section $3.$
@@
Top 1 accuracy 63.2 71.4 61.2 68.5 64.7 72.1 63.8 71.8
Figure 9: Adapting pre-trained ViTs into CBSA style. We finetune both the adapted models and the original ViTs on ImageNet-1K for 50 epochs, and report the top-1 accuracy on the validation set. CBSA leverages the three distinct projection matrices inherited from the pretrained MHSA to calculate query, key, and value, instead of using a single projection matrix as in (8). CBSA refers to CBSA without the contraction step, which essentially becomes Agent Attention.

### v3

@@ 3.2 Contract-and-Broadcast Self-Attention
Figure 3: Different attention mechanisms. $\top$ stands for the matrix transpose. For simplicity, the agent bias in the diagram of Agent Attention [39] are omitted, and the diagram of our CBSA is also simplified for comparison.
3.3 CBSA as a unified attention formula
In this subsection, we explore the potential of CBSA to serve as a unified formula for different attention mechanisms. Unlike recent work [40, 41], CBSA encompasses a broader spectrum of mechanisms (see Fig. 3) in an interpretable and mathematically grounded manner.
Our analysis reveals that, CBSA (7) can derive multiple variants corresponding to existing attention mechanisms by varying the choice of representatives. In these variants, the distinct initialization, extraction, contraction, and broadcast steps of CBSA may not be explicitly observed, as some of them are simplified or fused together due to the specific number and structure of representatives, or engineering concerns.
Softmax attention variant. Obviously, the input tokens themselves satisfy the constraints in (4), and thus can be directly used as the representatives, i.e., $\boldsymbol{Q}=\boldsymbol{Z}$ and $m=N$. In this case, we call the input tokens are self-expressed [42], where data samples are linearly represented over a dictionary composed of themselves. Then we can take a trivial solution for the regularization term where all coefficient matrices are identity matrices. Substituting them into (8) yields the following operator, known as MSSA in the white-box transformer [23]:
@@ 3.3 CBSA as a unified attention formula
Linear attention variant. To analyze the case of orthogonal representatives, we start with a canonical choice: the principal directions of input tokens. We thus perform singular value decomposition (SVD) within each subspace:
@@
where $\boldsymbol{L}_{k}\in\mathcal{O}(p)$, $\boldsymbol{R}_{k}\in\mathcal{O}(N,p)$, $\boldsymbol{\Sigma}_{k}$ is a $p\times p$ diagonal matrix of singular values, and the columns of $\boldsymbol{L}_{k}$ are known as the principal directions. Then, by right multiplying both sides by $\boldsymbol{R}_{k}$, we have:
$\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}\boldsymbol{R}_{k}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\boldsymbol{R}_{k}^{\top}\boldsymbol{R}_{k}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k},\forall\,k\in[K].$ (11)
By comparing (11) with the way to form the representatives, i.e., $\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}\boldsymbol{A}_{k}$, we let $\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}$ and $\boldsymbol{A}_{k}=\boldsymbol{R}_{k}$. Substituting them into (7) leads to the following operator:
@@
where $\mathcal{F}$ is a function defined on the spectrum of a positive semi-definite matrix and applies $f(\lambda_{i})=\nicefrac{{\epsilon^{2}}}{{(\epsilon^{2}+\lambda_{i})}}$ to each eigenvalues $\{\lambda_{i}\}_{i=1}^{p}$ of the covariance matrix. [^6] As $f$ is monotonically decreasing, the effect of (12) is to preserve the principle directions (i.e., representatives) with large variance while suppressing the other directions with vanishing variance [31]. This operator highly resembles the linear attention [21], due to that it also factorizes the $N\times N$ attention matrix and multiplies the key and value first to linearize the computational complexity. [^7] In fact, (12) is less expressive than black-box linear attention, because $(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})^{\top}$ is symmetric but $(\boldsymbol{W}_{\text{\!\!value}}\boldsymbol{Z})(\boldsymbol{W}_{\text{\!\!key}}\boldsymbol{Z})^{\top}$ is generally not. In Appendix B, we prove that a similar result holds for any set of orthogonal representatives.
@@
Channel attention variant. Assuming that the basis vectors of $\boldsymbol{U}_{k}$ are the principal directions for any set of input tokens (which is impossible but simplifies the computation), the directions of the representatives can be fixed along these basis vectors, i.e., $\boldsymbol{L}_{k}=\mathbf{I}_{p}$, thereby being orthogonal and input-agnostic (fixed). Then, (12) is simplified to:
@@
$\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{D}_{k}\boldsymbol{U}_{k}^{\top}\boldsymbol{Z},\ \ \text{where}\ \ \ \boldsymbol{D}_{k}\doteq\operatorname{Diag}\left([f\big((\boldsymbol{u}_{ki}^{\top}\boldsymbol{Z})(\boldsymbol{u}_{ki}^{\top}\boldsymbol{Z})^{\top}\big)]_{i=1}^{p}\right),$ (13)
@@
which basically recovers TSSA [24]. In (13), the feature channels are adaptively scaled according to their second moments of token projections, while channel attention [25, 26] typical employs an MLP to predict the channel-wise scaling factors.
Agent attention variant. Agent Attention is basically a variant of CBSA with the contraction step removed [^8] Since the query, key, and value are identical in CBSA, the broadcast term can be obtained directly from the extraction step, eliminating the separate computation branch in Agent Attention. which can be perceived in Fig. 3. Although this removal appears to confine the token-mixing ability, it is compensated by the pooling-based initialization, which is also a token mixer [43].
To gain some intuition of the gap in expressive capacity among the aforementioned mechanisms, we illustrate their compression patterns in the right panel of Fig. 1. The channel attention variant is restricted to compressing input tokens along fixed axes parameterized by $\boldsymbol{U}_{[K]}$, whereas the linear attention variant compresses them along principal directions that are dynamically determined by the input. We argue that such a dynamism is crucial for in-context learning [7] and for mitigating superposition [44]. In contrast, softmax attention exhibits much greater flexibility, as it manipulates each token independently. Actually, its compression can be viewed as operating in an $N$-dimensional space, rather than in the $d$-dimensional feature space [45]. Our proposed CBSA aims to approximate the behavior of softmax attention while significantly reducing computational cost.
@@
The above findings can also be interpreted from a dictionary learning perspective, where the representatives correspond to the atoms of a dictionary. When the representatives are orthogonal and fixed, they form a complete dictionary; when they are orthogonal yet input-dependent, they resemble a submatrix of an overcomplete dictionary as in compressed sensing [46]. When the input tokens themselves serve as representatives, they constitute a self-expressive dictionary [42].
@@ 4 Experiments
In this section, we evaluate the interpretability and the efficiency of the proposed CBSA and the CBTs built upon CBSA. As natural images often lie on low-dimensional subspaces [47, 48], we focus on classical visual tasks such as image classification and semantic segmentation, where higher resolutions generally lead to better accuracy [49, 50].
Baseline and training configuration. We compare our CBSA to the vanilla softmax attention [38, 1], and interpretable attention mechanisms based on MCR2, e.g., CRATE [23], ToST [24] and DEPICT [36]. Table 2 summarizes the baselines with brief descriptions. The results in gray are cited directly from the corresponding papers; whereas the others are reproduced under varied settings for fair comparisons. By default, the training configuration follows the baselines, with detailed information provided in Appendix C.
Implementation detail. The projection back to the ambient space, which should theoretically be a left multiplication by $\boldsymbol{U}_{k}$, is over-parameterized with an independently learnable matrix. This strategy is also adopted in MSSA [23] and TSSA [24], and its effect has been analyzed in [37]. In short, although this relaxation compromises the theoretical rigour, it is crucial for achieving better accuracy. In addition, the step sizes $\eta$ in (5) and $\kappa$ in (6) are implemented as learnable parameters without constraining their sign. This enables the model to adaptively switch between compression and decompression as needed. The PyTorch implementation is provided in Appendix D.
Table 2: Summary of baselines. Note that these methods are not limited to the tasks listed here, our descriptions only indicate their usages in the experiments of this paper.
Methods Attention Mechanism Complexity Interpretable Tasks
@@
DEPICT [36] MSSA quadratic semantic segmentation
@@ 4.1 Advantages enabled by interpretability
In this subsection, we show superior advantages of CBSA over black-box attention mechanisms. The most essential aspect of our CBSA is its interpretability, which induces other desirable properties such as robustness and emergent segmentation.
@@
Figure 4: Compact and structured representation. There are 10 classes indicated by colors. Points in each class are generated by sampling from a one-dimensional subspace and are then perturbed by adding noise.
Compact and structured representation. Similar to MCR2 [22], our optimization objective (4) aims to learn compact and structured representation by compressing input tokens towards low-dimensional subspaces. To confirm whether its iterative gradient steps can actually achieve this goal, we iterates the linear attention variant of CBSA (12) on synthetic data, where image tokens are modeled as the points in a three-dimensional space $\mathbb{R}^{3}$. Specifically, it is conducted on each class in the ambient space (thus being parameter-free) with a forward-only manner. As shown in Fig. 4, the representation ultimately admits a union of well-separated one-dimensional subspaces after $1024$ iterations.
@@
Figure 5: Evaluation on compression effect. We measure the compression term of (2) as a function of layers.
@@
Compressing all by contracting a few. To confirm our interpretation that CBSA compresses all input tokens by contracting a few representatives, we check that: whether the input tokens are indeed compressed; if so, whether the compression is driven by contracting the representatives. We measure the compression term of (2) as well as its normalized variant [^9] That is, the input tokens are normalized to unit vectors before measuring their coding rate. in Fig. 5. This normalized coding rate is invariant to in-place scaling and depends on the angles between input tokens. We observe two pieces of supporting evidences: a) the more compact the ultimate representation is, i.e., the compression term measured in the last layer is lower, the better the model performs on ImageNet-1K (see Table 3); [^10] Note that this correlation does not hold for the normalized coding rate, indicating that compression in magnitude is a critical aspect. b) the latter half of the layers exhibit consecutive compression, in nearly all models. [^11] We regard the decompression phenomenon in the early layers as a “known unknown” requiring further investigation. Then, in Fig. 6, we measure the reduced coding rate of the input tokens and representatives, respectively, after they are processed by CBSA within each subspace. We observe that the two kinds of reduced coding rates show highly similar trends across most subspaces. [^12] We hypothesis that the higher coding rate of the representatives makes them act as “scalpels” in order to give a “surgery” on the input tokens.
@@
Figure 6: Comparison on reduced coding rates. The reduction with respect to input tokens is computed between the input and their compressed counterparts, whereas the reduction with respect to representatives is computed between the extracted and the contracted representatives. We evaluate the reduced coding rate for both cases across all heads of the model. Results for CBT-T/S are shown here, and those for CBT-B/L are provided in Fig. 10. Note that $\kappa$ is fixed to 11 to exclude the decompression cases observed in Fig. 5.
@@
Figure 7: Visualization of [CLS] attention maps. We empirically estimate the full attention matrix by $\boldsymbol{A}_{k}\boldsymbol{A}_{k}^{\top}\in\mathbb{R}^{N\times N}$ and visualize the attention maps of the [CLS] token from the early, middle, and late layers of CBT, CRATE, and their hybrid model, respectively.
@@
Emergent segmentation properties. It has been reported that segmentation properties emerge in CRATE with merely standard supervised classification training owing to MSSA [52]. Compared to CRATE, our CBT attends to more semantically meaningful regions in the early layers, but the segmentation properties fail to persist in subsequent layers, as shown in Fig. 7. To address this phenomenon, we construct a hybrid model, termed CRATE+CBT, where the first half of the attention layers employ MSSA and the latter half employs CBSA. In this hybrid model, we observed that the segmentation properties not only emerged in the very first layer, but are also progressively enhanced in the following layers, rather than fading as in CBT. Qualitative results supporting this conclusion can be found in Appendix C.
@@
Figure 8: Semantic segmentation on ADE20K. All results are evaluated on the ADE20K validation set. CBSA indicates that the CBSA layers are implemented rigorously, i.e., without the overparameterization trick. Middle and Right: Random Gaussian noises with zero mean and standard deviation $\sigma$ are independently added to each parameter of the attention layers in the decoder.
@@
Robustness against parameter perturbation. As the attention heads of CBSA are modeled as low-dimensional subspaces, perturbing their projection matrices (i.e., subspace bases) with relatively small noise does not significantly alter the subspaces they span [36]. Consequently, as shown in the right two panels of Fig. 8, CBSA is extremely robust against parameter perturbation, whereas the black-box method (i.e., Segmenter [51]) collapses under the same perturbation.
@@
Table 3: Classification on ImageNet-1k. Models are trained on images of resolution $224\times 224$ with a patch size of $16\times 16$ where CBT and ViT use convolutional embedding layers but CRATE uses linear embeddings.
@@
Table 4: Fair comparisons on ImageNet-1K. All models employ convolutional embedding layers and ISTA feedforward blocks. The only difference lies in the attention mechanism, where Agent-T and Agent-S  [39] are implemented as in CBSA but without the contraction step (as derived in Section 3.3).
@@
Top-1 accuracy 63.2 71.4 61.2 68.5 64.7 72.1 63.8 71.8
Figure 9: Adapting pre-trained ViTs into CBSA style. We finetune both the adapted models and the original ViTs on ImageNet-1k for 50 epochs, and report the top-1 accuracy on the validation set. CBSA leverages the three distinct projection matrices inherited from the pretrained MHSA to calculate the query, key, and value matrices, instead of using a single projection matrix as in (8). CBSA refers to the CBSA without the contraction step, which is essentially an Agent Attention.